# Provably Safe Model Updates

### v1

@@ IV-C Choice of Specification Function
Effectively applying alternating gradient descent-ascent algorithms and solving Equation 4 imposes specific requirements on the specification function $\phi$. Most notably, any specification function must be differentiable. While a convenient choice for $\phi$ is accuracy, due to its interpretability, we cannot directly use accuracy since it is non-differentiable. Thus, our experiments rely on using soft accuracy $\phi_{\mathrm{soft}}$ as a differentiable surrogate:
$\phi_{\mathrm{soft}}(\theta)=\frac{1}{N}\sum_{k=0}^{N}\sum_{c=1}^{C}y_{k,c}\text{Softmax}(f^{\theta}(x_{k}))_{c}$
@@ IV-D Early stopping and Checkpointing
To mitigate this effect our implementation saves the current LID periodically during the computation and terminates the LID optimization before convergence. This produces a set $\mathcal{O}$ of domain parameters each corresponding to a LID. We then issue a certificate of the minimum accuracy for each LID in $\mathcal{O}$. When training the next task, we must project the model parameters (the result of proposed updates) $\theta$ onto the set of LIDs $\mathcal{O}$. Several strategies could be used to choose the LID $o^{*}$ on which we project $\theta$. An intuitive extension of the general case above is projecting onto the closest LID using:
$o^{*}=\arg\min_{o\in\mathcal{O}}\|\Pi(\theta,o)-\theta\|_{2}^{2}$
where $\Pi$ is the projection operator defined as:
$\Pi(\theta,o)_{i}=\mathrm{Clamp}(\theta_{i},\theta^{L}_{i},\theta^{U}_{i}),\qquad o=[\theta^{L},\theta^{U}]$
Another strategy is to choose the LID $o^{*}$ that results in the best immediate performance on our current batch. Letting $\mathcal{B}:=\{(x^{(i)},y^{(i)}\}_{i=1}^{B}$ be a batch of data the loss :
$o^{*}=\arg\min_{o\in\mathcal{O}}\mathcal{L}(\Pi(\theta,o),\mathcal{B})$
@@
The introduction of checkpointing forced an adaptation in our overall algorithm to allow for multiple LIDs for a single model. This can be combined with aforementioned biasing methods to compute distinctly biased LIDs that can be adapted to different types of downstream tasks. Having LIDs that are expanded along separate dimensions of the parameter space, lends a significant amount of flexibility to our algorithm that predecessors like InterContiNet did not exploit [wolczyk2022continual].
@@
Figure 4: Performance on the Split-CIFAR10 [CIFAR] dataset under three continual learning scenarios. The top panel shows the average accuracy over all seen tasks without replay data, while the bottom panel the average performance over seen tasks, while having varying amounts of buffer data available. The dashed line and hashed bars display the certificates returned by our method. Error bars illustrate the 80% confidence interval over 15 runs.
@@ IV-E Algorithm Discussion
Despite its computational efficiency and differentiability, the IBP approximation we leverage is known to exhibit increasing looseness with network depth and width [katz2017reluplex, sosnin2024certified]. Thus, while arbitrarily large networks can be considered, parameter intervals over the early layers of the network contribute disproportionately to the overall interval bounds on the outputs [wicker2020probabilistic]. As a result, our LIDs likely exhibit larger intervals for the later layers of the network compared to the earlier layers. In sufficiently deep networks, this effect can lead to the early layers becoming effectively ‘frozen,’ with little (or no) flexibility for future adaptation, while the final layers, subject to less over-approximation, remain relatively less constrained. This difficulty means that ensuring useful guarantees for large models requires careful design and hyper-parameter choices; we provide further systematic discussion of this phenomenon in Appendix -C.
@@
In this section we discuss some of the potential future applications enabled by a framework that enables a priori safe update guarantees. Most directly, the requirement for an a priori safe update mechanism has been put forth as a regulatory requirement for adjusting models in some critical domains. For instance, the U.S. Food and Drug Administration (FDA) outlined guidelines that enable pre-planned model updates, but only if they can be demonstrated to be safe a priori [us2024predetermined]. Furthermore, regulatory attention has also been attracted by the wide-spread use of foundation models (e.g. large language models). Regulations surrounding foundation models may require persistent safety mechanisms, especially given that models must respond dynamically to emerging vulnerabilities. While a valid response would be model de-provisioning (taking the model offline), using known-safe model updates is preferable for availability-critical systems. In the context of updating provided or open-source large language models, a potential use case of our method could be guaranteeing the preservation of learned behavior (e.g. alignment) for down-stream fine-tuners. As we have discussed, if not done carefully, fine-tuning can rapidly erode alignment properties [anderljung2023frontier, qi2024evaluating]. Thus, allowing third-party users to safely tune a model while the publishing entity can ensure, a priori, that this individualized updating procedure does not yield a model that violates originally aligned behavior. Such a guarantee may lead to robust terms-of-use.
@@ V-A Continual Learning Experiments
Baselines We compare our method against the following established continual learning methods: (1) EWC [kirkpatrick2017overcoming]: a regularization approach that selectively weights important parameters based on Fisher information, (2) LwF [li2017learning]: a distillation-based method that preserves functional behavior by using current task samples, (3) SGD: sequential training across tasks without explicit mitigation of catastrophic forgetting and (4) InterContiNet (ICN): a weight-space-constraint based approach to avoiding catastrophic forgetting [wolczyk2022continual]. To benchmark our buffer extension, we observe our method compared to its standard, zero-buffer implementation and Averaged GEM (A-GEM) [chaudhry2019efficientlifelonglearningagem]. For each baseline we consider the main incremental learning protocols outlined in [van2022three]:
@@
We divide the standard MNIST dataset [lecun1998gradient] into five sequential contexts, each containing an odd and an even number. Digit allocation across contexts was randomized for each experimental run, with reported performance metrics averaged across 100 independent runs. Figure 3 presents our method’s performance across the three Split-MNIST scenarios. We train a feedforward neural network with 2 convolutional layers (each with 8 channels and a $5\times 5$ kernel), followed by a dense layer. In Task-IL, all continual learning approaches effectively acquire new tasks while preserving performance on previous ones. However, for the more challenging Class-IL and Domain-IL settings, average accuracy deteriorates as task count increases. Impressively, our approach demonstrates competitive performance in the Class-IL setting, while simultaneously providing theoretical guarantees (dashed line) that bound the minimum accuracy under any future learning. In Domain-IL, our method maintains competitive average performance compared to baseline approaches while exhibiting improved retention of first-task performance. Moreover, our method observes more flexibility in its constraints, whereas ICN is highly prone to constraint collapse, as evident by the depicted tightness of the bounds. Additionally, using buffer data can drastically increase the performance of our method with a large buffer outperforming even A-GEM in the most difficult continual learning setting, CIL, and strictly improving performance in Domain-IL. Task-IL does not show the same trend as the setting is too simple and allows for enough model flexibility for even the zero buffer case to extract consistently high performance.
@@
We now turn to the Split-CIFAR10 dataset and consider the training and subsequent fine-tuning of a pre-trained ResNet18 model. The entire model is first trained on the initial task, after which subsequent tasks are learned by fine-tuning only the final linear layer of the network. The 10 classes are randomly allocated to 5 tasks, following the three incremental learning settings described above. In Figure 4, we observe similar overall trends to Split-MNIST; specifically, our method is consistently competitive with baseline approaches, while also offering formal guarantees on forgetting performance. In Domain-IL, we observe consistently better accuracy than baselines though our worst-case guarantees are below the performance of baseline methods (which is generally to be expected). Class incremental learning remains extremely challenging for all methods–where we consider only preserving task 1 performance, our method significantly out-performs baselines.
@@ V-B1 Foundation Model Performance on Hate-Speech Classification
We evaluate our approach on the binary task of Hate-Speech Classification (non-hate, hate labels) for the following datasets: Jigsaw [jigsaw-toxic-comment-classification-challenge], Twitter Hate Speech [davidson2017automated], Civil Comments [civil_comments], Hatemoji [hatemoji], Sbic [sbic] and Hatecheck [hatecheck]. For that purpose we use the frozen embeddings of the following foundation models that at the time of writing have top performance on existing benchmarks : 3-large (OpenAI) [openai2024embedding], MXBAI (Mistral) [choi2024linq], M2-BERT (Together) [together2024m2bert], GTE-Large (Alibaba) [ye2022complementary], Voyage-3 (Voyage) [voyage2024voyage3].
@@ V-B2 Foundation Model Performance on Multilingual Sentiment Analysis
Following the same procedure as above, we additionally evaluate our approach on the multilingual classification task of sentiment analysis (negative, neutral, positive) for the following multilingual dataset : XLM-T [xlmt] which contains 8 splits corresponding to 8 different languages
We provide a subset of the results in Figure 7. The dotted gray line represents the baseline performance of Voyage-3 using EWC [kirkpatrick2017overcoming]. When looking at the performances on Gemini, Voyage-AI and 3-large, we observe that our method achieves performances comparable to the EWC baseline and even occasionally outperforms it. Additionally, our method certifies non-trivial bounds on the original task macro-accuracy. These results provide the first empirical evidence that formal forgetting/alignment-preserving guarantees can be achieved in real-world LLM-based NLP pipelines, without sacrificing utility on the new task. The certified safety margins, though conservative, are large enough to ensure that updates remain practically useful.
@@ V-C LID Biasing Experiments
In this section we explore ablations on our proposed LID biasing approaches by considering the Split-MNIST CIL setting. We keep the convolutional architecture described in prior experiments and consider a non-biased LID computation as a baseline. We explore biasing the optimization using the Fisher information matrix [aich2021elasticweightconsolidationewc] and synaptic intelligence [zenke2017synaptic_intelligence] to obtain weight importance metrics by training our model on a batch with either lookahead data or the original task data. All experiments in this section consider pruning rather than regularization.
Pruning Prop. 0.5 0.7 0.8 0.825 0.85 0.9 0.95
With Lookahead
Pruning Prop. 0.5 0.3 0.2 0.175 0.15 0.1 0.05
Standard (No Buffer)

### v2

@@ IV-C Choice of Specification Function
Effectively applying alternating gradient descent-ascent algorithms and solving Equation  (4) imposes specific requirements on the specification function $\phi$. Most notably, any specification function must be differentiable. While a convenient choice for $\phi$ is accuracy, due to its interpretability, we cannot directly use accuracy since it is non-differentiable. Thus, our experiments rely on using soft accuracy $\phi_{\mathrm{soft}}$ as a differentiable surrogate:
$\phi_{\mathrm{soft}}(\theta)=\frac{1}{N}\sum_{k=0}^{N}\sum_{c=1}^{C}y_{k,c}\text{Softmax}(f^{\theta}(x_{k}))_{c},$
@@ IV-D Early Stopping and Checkpointing
To mitigate this effect our implementation saves the current LID periodically during the computation and terminates the LID optimization before convergence. This produces a set $\mathcal{O}$ of domain parameters each corresponding to a LID. We then issue a certificate of the minimum accuracy for each LID in $\mathcal{O}$. When training the next task, we must project the model parameters (the result of proposed updates) $\theta$ onto the set of LIDs $\mathcal{O}$. Several strategies could be used to choose the LID $o^{*}$ on which we project $\theta$. An intuitive extension of the general case above is projecting onto the closest LID using
$o^{*}=\arg\min_{o\in\mathcal{O}}\|\Pi(\theta,o)-\theta\|_{2}^{2},$
where $\Pi$ is the projection operator defined as
$\Pi(\theta,o)_{i}=\mathrm{Clamp}(\theta_{i},\theta^{L}_{i},\theta^{U}_{i}),\qquad o=[\theta^{L},\theta^{U}].$
Another strategy is to choose the LID $o^{*}$ that results in the best immediate performance on our current batch. More specifically, letting $\mathcal{B}:=\{(x^{(i)},y^{(i)}\}_{i=1}^{B}$ be a batch of data we have that
$o^{*}=\arg\min_{o\in\mathcal{O}}\mathcal{L}(\Pi(\theta,o),\mathcal{B}),$
@@
The introduction of checkpointing forced an adaptation in our overall algorithm to allow for multiple LIDs for a single model. This can be combined with aforementioned biasing methods to compute distinctly biased LIDs that can be adapted to different types of downstream tasks. Having LIDs that are expanded along separate dimensions of the parameter space, lends a significant amount of flexibility to our algorithm that predecessors like InterContiNet did not exploit [58].
@@
Figure 4: Performance on the Split-CIFAR10 [25] dataset under three continual learning scenarios. The top panel shows the average accuracy over all seen tasks without replay data, while the bottom panel the average performance over seen tasks, while having varying amounts of buffer data available. The dashed line and hashed bars display the certificates returned by our method. Error bars illustrate the 80% confidence interval over 15 runs.
@@ IV-E Algorithm Discussion
Despite its computational efficiency and differentiability, the IBP approximation we leverage is known to exhibit increasing looseness with network depth and width [22, 44]. Thus, while arbitrarily large networks can be considered, parameter intervals over the early layers of the network contribute disproportionately to the overall interval bounds on the outputs [53]. As a result, our LIDs likely exhibit larger intervals for the later layers of the network compared to the earlier layers. In sufficiently deep networks, this effect can lead to the early layers becoming effectively ‘frozen,’ with little (or no) flexibility for future adaptation, while the final layers, subject to less over-approximation, remain relatively less constrained. This difficulty means that ensuring useful guarantees for large models requires careful design and hyper-parameter choices; we provide further systematic discussion of this phenomenon in Appendix -C.
@@
In this section we discuss some of the potential future applications enabled by a framework that enables a priori safe update guarantees. Most directly, the requirement for an a priori safe update mechanism has been put forth as a regulatory requirement for adjusting models in some critical domains. For instance, the U.S. Food and Drug Administration (FDA) outlined guidelines that enable pre-planned model updates, but only if they can be demonstrated to be safe a priori [47]. Furthermore, regulatory attention has also been attracted by the wide-spread use of foundation models (e.g. large language models). Regulations surrounding foundation models may require persistent safety mechanisms, especially given that models must respond dynamically to emerging vulnerabilities. While a valid response would be model de-provisioning (taking the model offline), using known-safe model updates is preferable for availability-critical systems. In the context of updating provided or open-source large language models, a potential use case of our method could be guaranteeing the preservation of learned behavior (e.g. alignment) for down-stream fine-tuners. As we have discussed, if not done carefully, fine-tuning can rapidly erode alignment properties [3, 34]. Thus, allowing third-party users to safely tune a model while the publishing entity can ensure, a priori, that this individualized updating procedure does not yield a model that violates originally aligned behavior. Such a guarantee may lead to robust terms-of-use.
@@ V-A Continual Learning Experiments
Baselines We compare our method against the following established continual learning methods: (1) EWC [24]: a regularization approach that selectively weights important parameters based on Fisher information, (2) LwF [29]: a distillation-based method that preserves functional behavior by using current task samples, (3) SGD: sequential training across tasks without explicit mitigation of catastrophic forgetting and (4) InterContiNet (ICN): a weight-space-constraint based approach to avoiding catastrophic forgetting [58]. To benchmark our buffer extension, we observe our method compared to its standard, zero-buffer implementation and Averaged GEM (A-GEM) [8]. For each baseline we consider the main incremental learning protocols outlined in [48]:
@@
We divide the standard MNIST dataset [27] into five sequential contexts, each containing an odd and an even number. Digit allocation across contexts was randomized for each experimental run, with reported performance metrics averaged across 100 independent runs. Figure 3 presents our method’s performance across the three Split-MNIST scenarios. We train a feedforward neural network with 2 convolutional layers (each with 8 channels and a $5\times 5$ kernel), followed by a dense layer. In Task-IL, all continual learning approaches effectively acquire new tasks while preserving performance on previous ones. However, for the more challenging Class-IL and Domain-IL settings, average accuracy deteriorates as task count increases. Impressively, our approach demonstrates competitive performance in the Class-IL setting, while simultaneously providing theoretical guarantees (dashed line) that bound the minimum accuracy under any future learning. In Domain-IL, our method maintains competitive average performance compared to baseline approaches while exhibiting improved retention of first-task performance. Moreover, our method observes more flexibility in its constraints, whereas ICN is highly prone to constraint collapse, as evident by the depicted tightness of the bounds.
Additionally, using buffer data can drastically increase the performance of our method with a large buffer outperforming even A-GEM in the most difficult continual learning setting, CIL, and strictly improving performance in Domain-IL. Task-IL does not show the same trend as the setting is too simple and allows for enough model flexibility for even the zero buffer case to extract consistently high performance.
@@
We now turn to the Split-CIFAR10 dataset and consider the training and subsequent fine-tuning of a pre-trained ResNet18 model. The entire model is first trained on the initial task, after which subsequent tasks are learned by fine-tuning only the final linear layer of the network. The 10 classes are randomly allocated to 5 tasks, following the three incremental learning settings described above. In Figure 4, we observe similar overall trends to Split-MNIST; specifically, our method is consistently competitive with baseline approaches, while also offering formal guarantees on forgetting performance. In Domain-IL, we observe consistently better accuracy than baselines though our worst-case guarantees are below the performance of baseline methods (which is generally to be expected). Class incremental learning remains extremely challenging for all methods—where we consider only preserving task 1 performance, our method significantly out-performs baselines.
@@ V-B1 Foundation Model Performance on Hate-Speech Classification
We evaluate our approach on the binary task of Hate-Speech Classification (non-hate, hate labels) for the following datasets: Jigsaw [10], Twitter Hate Speech [13], Civil Comments [7], Hatemoji [23], Sbic [42] and Hatecheck [39]. For that purpose we use the frozen embeddings of the following foundation models that at the time of writing have top performance on existing benchmarks : 3-large (OpenAI) [33], MXBAI (Mistral) [9], M2-BERT (Together) [11], GTE-Large (Alibaba) [60], Voyage-3 (Voyage) [1].
@@
With Lookahead Buffer
Pruning Prop. 0.5 0.7 0.8 0.825 0.85 0.9 0.95
No Lookahead Buffer (Standard)
Pruning Prop. 0.5 0.3 0.2 0.175 0.15 0.1 0.05
@@ V-B2 Foundation Model Performance on Multilingual Sentiment Analysis
Following the same procedure as above, we additionally evaluate our approach on the multilingual classification task of sentiment analysis (negative, neutral, positive) for the following multilingual dataset : XLM-T [4] which contains 8 splits corresponding to 8 different languages.
We provide a subset of the results in Figure 6. The dotted gray line represents the baseline performance of Voyage-3 using EWC [24]. When looking at the performances on Gemini, Voyage-AI and 3-large, we observe that our method achieves performances comparable to the EWC baseline and even occasionally outperforms it. Additionally, our method certifies non-trivial bounds on the original task macro-accuracy. These results provide the first empirical evidence that formal forgetting/alignment-preserving guarantees can be achieved in real-world LLM-based NLP pipelines, without sacrificing utility on the new task. The certified safety margins, though conservative, are large enough to ensure that updates remain practically useful.
@@ V-C LID Biasing Experiments
In this section we explore ablations on our proposed LID biasing approaches by considering the Split-MNIST CIL setting. We keep the convolutional architecture described in prior experiments and consider a non-biased LID computation as a baseline. We explore biasing the optimization using the Fisher information matrix [2] and synaptic intelligence [62] to obtain weight importance metrics by training our model on a batch with either lookahead data or the original task data. All experiments in this section consider pruning rather than regularization.